# A SAT-centered XAI method for Deep Learning based Video Understanding

### v1

@@ 1 Introduction and Background
SAT solvers have demonstrated their utility across various domains within artificial intelligence, largely due to their proficiency in managing intricate logical constraints. One notable application lies in hardware and software verification, where SAT solvers play a crucial role in validating the correctness of designs against formal specifications [45, 73]. For instance, in electronic design automation (EDA), SAT solvers are essential for formal equivalence checking, ensuring that different representations of a circuit design are logically consistent [22, 25]. They are also utilized in bounded model checking to verify that a system adheres to specific properties within a defined number of steps [33].
Another significant application area is automated planning and scheduling. Planning problems, such as determining a sequence of actions to achieve a particular goal, can be formulated as SAT problems. In this formulation, Boolean variables represent the state of the world at different time points, and clauses capture the preconditions and effects of actions [35, 36]. SAT solvers can then identify a satisfying assignment that corresponds to a valid plan. Similarly, scheduling problems, like allocating resources to tasks over time, can also be modeled and solved using SAT solvers [46, 39].
SAT solvers have also found applications in other AI tasks, including constraint satisfaction problems, diagnosis, and even certain aspects of machine learning. For example, the SATzilla portfolio SAT solver employs machine learning techniques to predict the runtime of different SAT solvers on a given instance and selects the most promising solver [54, 53]. The success of SAT solvers in addressing computationally challenging problems across diverse AI domains underscores their versatility and potential as a fundamental tool in the field [79, 78].
@@ 1.1 Formal Explainable Artificial Intelligence (XAI): Theoretical Underpinnings
Formal Explainable Artificial Intelligence (XAI) represents a shift in the pursuit of explainable AI, moving towards approaches rooted in formal logic and automated reasoning, rather than relying on heuristic and often post-hoc methods [71, 14]. The central concept of formal XAI is to provide explanations for the behavior and decisions of AI systems, especially complex machine learning models, using rigorous logic-based definitions that can be computed and verified using automated reasoning techniques [45, 72]. This approach seeks to overcome the limitations of many existing XAI methods that lack formal guarantees and can sometimes produce misleading information [47, 70].
@@ 1.1.1 Logic-Based Approaches to Explainability
Logic is fundamental in formal XAI, providing the basis for representing explanations and reasoning about the behavior of AI systems [65]. Symbolic AI, an earlier branch of AI, heavily utilized high-level symbolic representations and logic-based reasoning [69]. Modern formal XAI builds upon these principles, employing various logical formalisms to capture the reasoning processes of machine learning models and to formulate explanations in a precise and verifiable manner [23, 18]. These logical approaches offer the potential for provably correct explanations and the ability to formally reason about properties of the AI system, such as fairness or bias [12, 27].
The use of logic in explainability often involves representing the classifier or other machine learning model as a logical formula [51]. The input to the model can also be represented as a conjunction of literals, and the model’s prediction as another literal. The classification process can then be viewed through the lens of logical entailment [52, 50]. Explanations, in this context, can take the form of logical rules or sets of conditions that are necessary or sufficient for a particular prediction [42].
@@ 1.1.2 Abductive and Contrastive Explanations
Formal XAI often differentiates between types of explanations, such as abductive and contrastive explanations [49]. Abductive explanations aim to answer ”Why?” questions by providing a set of reasons or conditions that, if true, would explain the observed outcome or prediction [57, 58]. For a classifier, an abductive explanation for a given instance and its predicted class might be a minimal subset of features that, when fixed to their values in the instance, guarantee the same prediction [59, 60].
Contrastive explanations, conversely, address ”Why not?” questions by identifying the minimal changes to the input that would lead to a different outcome [8]. For example, a contrastive explanation might highlight which features need to be altered and by how much to change the prediction of a classifier from one class to another. Both abductive and contrastive explanations provide valuable and complementary perspectives on the model’s reasoning process, allowing for a more comprehensive understanding of its behavior [3, 5].
@@ 1.1.3 Formal Verification of AI Systems
Formal verification techniques are crucial for ensuring the correctness and reliability of AI systems, including the explanations they provide [9, 56]. Formal verification involves employing mathematical methods to prove that a system meets its predefined specifications or requirements [16, 26]. Techniques such as model checking, automated theorem proving, and deductive verification can be applied to analyze AI algorithms and their implementations to ensure they behave as expected under all possible conditions [30, 32].
In the context of explainability, formal verification can be used to assess the soundness and completeness of the generated explanations [67, 66]. Soundness would imply that if the explanation claims a certain set of factors led to a prediction, then those factors indeed guarantee that prediction according to the model’s logic. Completeness would mean that if a valid explanation exists within the defined framework, the verification process would be able to find it. While formal verification offers a high degree of rigor, it can be challenging to apply to very complex AI systems, like deep neural networks, due to their intricate nature and vast state spaces [55, 74, 1]. Nevertheless, it remains a crucial aspect of ensuring the trustworthiness and reliability of both AI systems and their explanations [48, 75, 77, 68, 15].
@@ 2.1 Combining SAT Solving and Formal XAI for Machine Learning Explanation
The integration of SAT solving techniques with the principles of formal XAI has emerged as a promising direction for generating rigorous explanations for machine learning models [73]. By leveraging the power of SAT solvers as automated reasoners, researchers have explored methods to extract logical explanations from various types of machine learning models, including classifiers [72].
@@ 2.1.1 Using SAT for Explaining Classifiers and Other ML Models
SAT solvers can be employed to explain the predictions of classifiers by formulating the classification process and the explanation query as a satisfiability problem [79]. For instance, to find an abductive explanation for a specific prediction, one could encode the classifier’s behavior as a set of logical clauses. Then, the input instance and the predicted class are also encoded. The explanation query could be formulated as finding a minimal subset of the input features such that, when their values are fixed, the resulting SAT problem (representing the classifier and the fixed features) still entails the predicted class [78]. If a satisfying assignment exists under these conditions, it indicates that the chosen subset of features is sufficient to guarantee the prediction. SAT solvers can then be used to find such minimal subsets, providing an explanation in terms of the most relevant input features [47, 70].
Similarly, contrastive explanations can be sought by encoding the condition that the model should predict a different class. The SAT solver can then identify the minimal changes to the input features that satisfy this condition, thus providing a ”why not?” explanation [13]. This approach has been applied to various interpretable machine learning models, such as decision trees and rule-based systems, as well as to more complex models like neural networks through appropriate encodings [54].
@@ 2.1.2 Logic-Based Rule Extraction from Machine Learning Models
Another way SAT solvers can contribute to explainability is through logic-based rule extraction from trained machine learning models [71]. The goal of rule extraction is to identify a set of human-readable logical rules that accurately capture the decision-making logic of the model. SAT solvers can be used in this process to find minimal sets of rules that are consistent with the model’s predictions on the training data [14]. For example, the model’s behavior on the training set can be encoded as a SAT problem, where variables represent features and their values, and clauses represent the model’s predictions. By finding satisfying assignments to this SAT problem under certain constraints (e.g., minimizing the number or complexity of the rules), one can extract a set of logical rules that effectively mimic the model’s behavior and can serve as explanations for its predictions [15].
These extracted rules can provide a more transparent and understandable representation of the model’s internal logic compared to the complex mathematical operations within the model itself. They can highlight the key features and their combinations that lead to specific predictions, making it easier for humans to understand and trust the model’s decisions [41, 27].
@@ 2.2 Explainable AI for Deep Learning Models
Explainable AI for deep learning models is a rapidly evolving field that explores various techniques to shed light on the decision-making processes of these complex architectures [62]. While many of these techniques are not formal in nature, understanding the current landscape is crucial for positioning the proposed SAT-based approach [65].
@@ 2.2.1 Post-hoc Explanation Techniques for Neural Networks
A significant portion of research in deep learning explainability focuses on post-hoc methods, which are applied to a trained model to understand its behavior without necessarily altering its architecture or training process [48]. Popular post-hoc techniques include:
  • Saliency Maps: These methods compute a relevance score for each input feature (e.g., pixels in an image, words in a text) indicating its importance for the model’s prediction [16]. Various techniques exist, such as gradient-based methods, which highlight the input regions that have the strongest influence on the output.
  • Attention Mechanisms: Many modern deep learning models, particularly in areas like natural language processing and video understanding, utilize attention mechanisms [26]. These mechanisms learn to weigh the importance of different parts of the input when making a prediction. The attention weights can often be interpreted as indicating which input elements the model focused on, providing a form of explanation.
  • LIME (Local Interpretable Model-agnostic Explanations): LIME works by approximating the behavior of the complex deep learning model locally around a specific prediction using a simpler, interpretable model (e.g., a linear model or a decision tree) [60]. By analyzing this local surrogate model, one can gain insights into which input features were most important for that particular prediction.
@@
These techniques provide valuable insights into the behavior of deep learning models, often offering intuitive visual or feature-based explanations. However, they often lack formal guarantees and can be sensitive to the specific implementation or parameters used [55, 74].
@@ 2.2.2 Limitations of Existing Methods in Providing Formal Guarantees
A major limitation of many existing XAI methods for deep learning is the absence of formal guarantees regarding the correctness or completeness of the explanations they provide [68]. While these methods can often highlight potentially relevant features or regions, they do not necessarily provide a logical justification for the model’s prediction in a way that can be formally verified. This lack of rigor can be problematic, especially in critical applications where trust and accountability are paramount. For instance, a saliency map might highlight a certain region in an image, but it does not formally prove that the model’s decision was solely based on that region or explain the underlying reasoning process. The potential for misleading information from informal XAI methods has been demonstrated, underscoring the need for more rigorous approaches [44].
@@ 2.3 Explainable AI in the Domain of Video Understanding
Explainable AI in the domain of video understanding presents unique challenges due to the spatio-temporal nature of video data and the complexity of deep learning models used for video analysis [58].
@@ 2.3.1 Current Techniques and Their Shortcomings
  • Spatio-Temporal Attention: Models that utilize attention mechanisms can provide insights into both the spatial regions and the temporal segments that the model deems important for its prediction [57]. Visualizing these attention weights across frames can offer a spatio-temporal explanation.
  • Gradient-Based Methods for Video: Gradient-based saliency methods can be extended to video by computing gradients with respect to both spatial and temporal inputs. This can highlight the video frames or regions within frames that are most influential for the prediction [8].
  • Temporal Perturbation Analysis: This involves perturbing or masking different temporal segments of the video and observing the impact on the model’s prediction. By identifying which segments are most critical, one can gain insights into the model’s temporal reasoning [3].
While these techniques offer some level of explanation for video understanding models, they often face limitations in providing comprehensive and formal explanations for video-level predictions. Explaining how long-term temporal relationships between events contribute to the final prediction remains a significant challenge [5, 9]. Furthermore, the high computational cost of analyzing and explaining complex video models can be a limiting factor [56, 4].
@@ 2.3.2 Challenges Specific to Explaining Spatio-Temporal Models
  • Attribution to Specific Regions and Segments: It is difficult to precisely attribute a prediction to specific spatial regions within individual frames and to specific temporal segments across the entire video. The model’s decision might be based on a complex interplay of both spatial and temporal features [43, 31].
  • Explaining Temporal Relationships: Understanding how the temporal relationships between different events or movements in a video contribute to the final prediction is particularly challenging. Many existing methods struggle to capture these complex temporal dependencies in their explanations [7, 11].
  • Computational Complexity: Analyzing and explaining deep learning models for video, which often involve processing large amounts of high-dimensional data, can be computationally very expensive. This limits the scalability and applicability of some explanation techniques [25, 22].
@@ 2.3.3 State-of-the-Art Deep Learning Models for Video Understanding and Their Explainability Bottlenecks
State-of-the-art deep learning models for video understanding often employ complex architectures such as 3D Convolutional Neural Networks (CNNs), recurrent neural networks (RNNs) like LSTMs and GRUs, and increasingly, Transformer-based models [10, 6]. These models can capture intricate spatio-temporal patterns but are often deeply layered and highly non-linear, making their internal representations and decision-making processes very difficult to interpret [63, 30].
For instance, 3D CNNs extend 2D convolutions to the temporal dimension, allowing them to learn spatio-temporal features directly from video data [61, 21]. However, understanding which specific spatio-temporal features are most important for a given prediction can be challenging [28, 29]. RNNs can model temporal dependencies, but the information is processed sequentially through hidden states, making it difficult to trace back the influence of specific past events on the current prediction [35, 32]. Transformer-based models, while powerful in capturing long-range dependencies, often have a large number of parameters and complex attention mechanisms that can be hard to interpret [36, 24].
The inherent complexity of these state-of-the-art models presents significant explainability bottlenecks. Their distributed representations and non-linear interactions make it hard to understand the causal relationships between the input video and the final prediction. This lack of transparency hinders trust, debugging, and further improvement of these powerful video understanding systems [20, 37, 38, 19, 2, 69, 34].
@@ 3 Proposed Method: Formal SAT-Based Explanation Model for Deep Video Understanding
This section introduces a novel method for providing formal, logic-based explanations for deep learning models applied to video understanding tasks. The proposed approach integrates SAT solving techniques with the principles of formal XAI to address the limitations of existing explanation methods in this domain [78].
@@ 3.1 Conceptual Framework: Integrating SAT Solving and Formal XAI for Video Analysis
The core idea of the proposed method is to leverage the rigorous reasoning capabilities of SAT solvers to explain the predictions made by deep learning models on video data [79]. This is achieved by creating a formal, logical representation of the deep learning model and the input video, and then formulating explanation queries as satisfiability problems within this logical framework. By solving these SAT problems, we can extract logical explanations for the model’s behavior [70].
@@
    Formal Encoding of the Deep Learning Model: A chosen deep learning model for video understanding is translated into a set of Boolean variables and logical clauses. This encoding captures the model’s architecture, parameters (weights and biases), and activation functions [47].
@@
    Formal Representation of Video Data: The input video is represented within the same logical framework. This might involve encoding features extracted from the video frames and the temporal relationships between them [13].
@@
    Formulation of Explanation Queries as SAT Problems: Specific explanation queries are translated into logical constraints that are added to the encoded model and video data. For example, a ”Why?” query might involve finding a minimal set of input features or internal activations that are necessary for the observed prediction. A ”Why not?” query could involve finding minimal changes to the input that would lead to a different prediction [54].
@@
    SAT Solving and Explanation Extraction: A SAT solver is then used to find a satisfying assignment to the resulting set of logical clauses. The solution provided by the SAT solver is interpreted back in the context of the original model and video data to yield a formal explanation [71, 14].
This framework aims to provide explanations that are not only interpretable but also formally verifiable, addressing the lack of rigor in many existing XAI techniques for deep learning in video understanding [15].
@@
Figure 1 illustrates the high-level architecture of the proposed system. It comprises a pre-trained deep learning model for video understanding, an encoding module that translates the model and the input video into a SAT problem, a query formulation module that defines the explanation query as logical constraints, a SAT solver, and an explanation extraction module that interprets the solver’s output [41, 27].

### v2

@@ 1 Introduction and Background
SAT solvers have demonstrated their utility across various domains within artificial intelligence, largely due to their proficiency in managing intricate logical constraints [2]. One notable application lies in hardware and software verification, where SAT solvers play a crucial role in validating the correctness of designs against formal specifications. For instance, in electronic design automation (EDA), SAT solvers are essential for formal equivalence checking, ensuring that different representations of a circuit design are logically consistent. They are also utilized in bounded model checking to verify that a system adheres to specific properties within a defined number of steps.
Another significant application area is automated planning and scheduling. Planning problems, such as determining a sequence of actions to achieve a particular goal, can be formulated as SAT problems. In this formulation, Boolean variables represent the state of the world at different time points, and clauses capture the preconditions and effects of actions. SAT solvers can then identify a satisfying assignment that corresponds to a valid plan. Similarly, scheduling problems, like allocating resources to tasks over time, can also be modeled and solved using SAT solvers.
SAT solvers have also found applications in other AI tasks, including constraint satisfaction problems, diagnosis, and even certain aspects of machine learning. For example, the SATzilla portfolio SAT solver employs machine learning techniques to predict the runtime of different SAT solvers on a given instance and selects the most promising solver. The success of SAT solvers in addressing computationally challenging problems across diverse AI domains underscores their versatility and potential as a fundamental tool in the field.
@@ 1.1 Formal Explainable Artificial Intelligence (XAI): Theoretical Underpinnings
Formal Explainable Artificial Intelligence (XAI) represents a shift in the pursuit of explainable AI [1], moving towards approaches rooted in formal logic and automated reasoning, rather than relying on heuristic and often post-hoc methods. The central concept of formal XAI is to provide explanations for the behavior and decisions of AI systems, especially complex machine learning models, using rigorous logic-based definitions that can be computed and verified using automated reasoning techniques. This approach seeks to overcome the limitations of many existing XAI methods that lack formal guarantees and can sometimes produce misleading information.
@@ 1.1.1 Logic-Based Approaches to Explainability
Logic is fundamental in formal XAI, providing the basis for representing explanations and reasoning about the behavior of AI systems. Symbolic AI, an earlier branch of AI, heavily utilized high-level symbolic representations and logic-based reasoning. Modern formal XAI builds upon these principles, employing various logical formalisms to capture the reasoning processes of machine learning models and to formulate explanations in a precise and verifiable manner. These logical approaches offer the potential for provably correct explanations and the ability to formally reason about properties of the AI system, such as fairness or bias.
The use of logic in explainability often involves representing the classifier or other machine learning model as a logical formula. The input to the model can also be represented as a conjunction of literals, and the model’s prediction as another literal. The classification process can then be viewed through the lens of logical entailment. Explanations, in this context, can take the form of logical rules or sets of conditions that are necessary or sufficient for a particular prediction.
@@ 1.1.2 Abductive and Contrastive Explanations
Formal XAI often differentiates between types of explanations, such as abductive and contrastive explanations. Abductive explanations aim to answer ”Why?” questions by providing a set of reasons or conditions that, if true, would explain the observed outcome or prediction. For a classifier, an abductive explanation for a given instance and its predicted class might be a minimal subset of features that, when fixed to their values in the instance, guarantee the same prediction.
Contrastive explanations, conversely, address ”Why not?” questions by identifying the minimal changes to the input that would lead to a different outcome. For example, a contrastive explanation might highlight which features need to be altered and by how much to change the prediction of a classifier from one class to another. Both abductive and contrastive explanations provide valuable and complementary perspectives on the model’s reasoning process, allowing for a more comprehensive understanding of its behavior.
@@ 1.1.3 Formal Verification of AI Systems
Formal verification techniques are crucial for ensuring the correctness and reliability of AI systems, including the explanations they provide. Formal verification involves employing mathematical methods to prove that a system meets its predefined specifications or requirements. Techniques such as model checking, automated theorem proving, and deductive verification can be applied to analyze AI algorithms and their implementations to ensure they behave as expected under all possible conditions.
In the context of explainability, formal verification can be used to assess the soundness and completeness of the generated explanations. Soundness would imply that if the explanation claims a certain set of factors led to a prediction, then those factors indeed guarantee that prediction according to the model’s logic. Completeness would mean that if a valid explanation exists within the defined framework, the verification process would be able to find it. While formal verification offers a high degree of rigor, it can be challenging to apply to very complex AI systems, like deep neural networks, due to their intricate nature and vast state spaces. Nevertheless, it remains a crucial aspect of ensuring the trustworthiness and reliability of both AI systems and their explanations.
@@ 2.1 Combining SAT Solving and Formal XAI for Machine Learning Explanation
The integration of SAT solving techniques with the principles of formal XAI has emerged as a promising direction for generating rigorous explanations for machine learning models. By leveraging the power of SAT solvers as automated reasoners, researchers have explored methods to extract logical explanations from various types of machine learning models, including classifiers.
@@ 2.1.1 Using SAT for Explaining Classifiers and Other ML Models
SAT solvers can be employed to explain the predictions of classifiers by formulating the classification process and the explanation query as a satisfiability problem. For instance, to find an abductive explanation for a specific prediction, one could encode the classifier’s behavior as a set of logical clauses. Then, the input instance and the predicted class are also encoded. The explanation query could be formulated as finding a minimal subset of the input features such that, when their values are fixed, the resulting SAT problem (representing the classifier and the fixed features) still entails the predicted class. If a satisfying assignment exists under these conditions, it indicates that the chosen subset of features is sufficient to guarantee the prediction. SAT solvers can then be used to find such minimal subsets, providing an explanation in terms of the most relevant input features.
Similarly, contrastive explanations can be sought by encoding the condition that the model should predict a different class. The SAT solver can then identify the minimal changes to the input features that satisfy this condition, thus providing a ”why not?” explanation. This approach has been applied to various interpretable machine learning models, such as decision trees and rule-based systems, as well as to more complex models like neural networks through appropriate encodings.
@@ 2.1.2 Logic-Based Rule Extraction from Machine Learning Models
Another way SAT solvers can contribute to explainability is through logic-based rule extraction from trained machine learning models. The goal of rule extraction is to identify a set of human-readable logical rules that accurately capture the decision-making logic of the model. SAT solvers can be used in this process to find minimal sets of rules that are consistent with the model’s predictions on the training data. For example, the model’s behavior on the training set can be encoded as a SAT problem, where variables represent features and their values, and clauses represent the model’s predictions. By finding satisfying assignments to this SAT problem under certain constraints (e.g., minimizing the number or complexity of the rules), one can extract a set of logical rules that effectively mimic the model’s behavior and can serve as explanations for its predictions.
These extracted rules can provide a more transparent and understandable representation of the model’s internal logic compared to the complex mathematical operations within the model itself. They can highlight the key features and their combinations that lead to specific predictions, making it easier for humans to understand and trust the model’s decisions.
@@ 2.2 Explainable AI for Deep Learning Models
Explainable AI for deep learning models is a rapidly evolving field that explores various techniques to shed light on the decision-making processes of these complex architectures. While many of these techniques are not formal in nature, understanding the current landscape is crucial for positioning the proposed SAT-based approach.
@@ 2.2.1 Post-hoc Explanation Techniques for Neural Networks
A significant portion of research in deep learning explainability focuses on post-hoc methods, which are applied to a trained model to understand its behavior without necessarily altering its architecture or training process. Popular post-hoc techniques include:
  • Saliency Maps: These methods compute a relevance score for each input feature (e.g., pixels in an image, words in a text) indicating its importance for the model’s prediction. Various techniques exist, such as gradient-based methods, which highlight the input regions that have the strongest influence on the output.
  • Attention Mechanisms: Many modern deep learning models, particularly in areas like natural language processing and video understanding, utilize attention mechanisms. These mechanisms learn to weigh the importance of different parts of the input when making a prediction. The attention weights can often be interpreted as indicating which input elements the model focused on, providing a form of explanation.
  • LIME (Local Interpretable Model-agnostic Explanations): LIME works by approximating the behavior of the complex deep learning model locally around a specific prediction using a simpler, interpretable model (e.g., a linear model or a decision tree). By analyzing this local surrogate model, one can gain insights into which input features were most important for that particular prediction.
@@
These techniques provide valuable insights into the behavior of deep learning models, often offering intuitive visual or feature-based explanations. However, they often lack formal guarantees and can be sensitive to the specific implementation or parameters used.
@@ 2.2.2 Limitations of Existing Methods in Providing Formal Guarantees
A major limitation of many existing XAI methods for deep learning is the absence of formal guarantees regarding the correctness or completeness of the explanations they provide. While these methods can often highlight potentially relevant features or regions, they do not necessarily provide a logical justification for the model’s prediction in a way that can be formally verified. This lack of rigor can be problematic, especially in critical applications where trust and accountability are paramount. For instance, a saliency map might highlight a certain region in an image, but it does not formally prove that the model’s decision was solely based on that region or explain the underlying reasoning process. The potential for misleading information from informal XAI methods has been demonstrated, underscoring the need for more rigorous approaches.
@@ 2.3 Explainable AI in the Domain of Video Understanding
Explainable AI in the domain of video understanding presents unique challenges due to the spatio-temporal nature of video data and the complexity of deep learning models used for video analysis.
@@ 2.3.1 Current Techniques and Their Shortcomings
  • Spatio-Temporal Attention: Models that utilize attention mechanisms can provide insights into both the spatial regions and the temporal segments that the model deems important for its prediction. Visualizing these attention weights across frames can offer a spatio-temporal explanation.
  • Gradient-Based Methods for Video: Gradient-based saliency methods can be extended to video by computing gradients with respect to both spatial and temporal inputs. This can highlight the video frames or regions within frames that are most influential for the prediction.
  • Temporal Perturbation Analysis: This involves perturbing or masking different temporal segments of the video and observing the impact on the model’s prediction. By identifying which segments are most critical, one can gain insights into the model’s temporal reasoning.
While these techniques offer some level of explanation for video understanding models, they often face limitations in providing comprehensive and formal explanations for video-level predictions. Explaining how long-term temporal relationships between events contribute to the final prediction remains a significant challenge. Furthermore, the high computational cost of analyzing and explaining complex video models can be a limiting factor.
@@ 2.3.2 Challenges Specific to Explaining Spatio-Temporal Models
  • Attribution to Specific Regions and Segments: It is difficult to precisely attribute a prediction to specific spatial regions within individual frames and to specific temporal segments across the entire video. The model’s decision might be based on a complex interplay of both spatial and temporal features.
  • Explaining Temporal Relationships: Understanding how the temporal relationships between different events or movements in a video contribute to the final prediction is particularly challenging. Many existing methods struggle to capture these complex temporal dependencies in their explanations.
  • Computational Complexity: Analyzing and explaining deep learning models for video, which often involve processing large amounts of high-dimensional data, can be computationally very expensive. This limits the scalability and applicability of some explanation techniques.
@@ 2.3.3 State-of-the-Art Deep Learning Models for Video Understanding and Their Explainability Bottlenecks
State-of-the-art deep learning models for video understanding often employ complex architectures such as 3D Convolutional Neural Networks (CNNs), recurrent neural networks (RNNs) like LSTMs and GRUs, and increasingly, Transformer-based models. These models can capture intricate spatio-temporal patterns but are often deeply layered and highly non-linear, making their internal representations and decision-making processes very difficult to interpret.
For instance, 3D CNNs extend 2D convolutions to the temporal dimension, allowing them to learn spatio-temporal features directly from video data. However, understanding which specific spatio-temporal features are most important for a given prediction can be challenging. RNNs can model temporal dependencies, but the information is processed sequentially through hidden states, making it difficult to trace back the influence of specific past events on the current prediction. Transformer-based models, while powerful in capturing long-range dependencies, often have a large number of parameters and complex attention mechanisms that can be hard to interpret.
The inherent complexity of these state-of-the-art models presents significant explainability bottlenecks. Their distributed representations and non-linear interactions make it hard to understand the causal relationships between the input video and the final prediction. This lack of transparency hinders trust, debugging, and further improvement of these powerful video understanding systems.
@@ 3 Proposed Method: Formal SAT-Based Explanation Model for Deep Video Understanding
This section introduces a novel method for providing formal, logic-based explanations for deep learning models applied to video understanding tasks. The proposed approach integrates SAT solving techniques with the principles of formal XAI to address the limitations of existing explanation methods in this domain.
@@ 3.1 Conceptual Framework: Integrating SAT Solving and Formal XAI for Video Analysis
The core idea of the proposed method is to leverage the rigorous reasoning capabilities of SAT solvers to explain the predictions made by deep learning models on video data. This is achieved by creating a formal, logical representation of the deep learning model and the input video, and then formulating explanation queries as satisfiability problems within this logical framework. By solving these SAT problems, we can extract logical explanations for the model’s behavior.
@@
    Formal Encoding of the Deep Learning Model: A chosen deep learning model for video understanding is translated into a set of Boolean variables and logical clauses. This encoding captures the model’s architecture, parameters (weights and biases), and activation functions.
@@
    Formal Representation of Video Data: The input video is represented within the same logical framework. This might involve encoding features extracted from the video frames and the temporal relationships between them.
@@
    Formulation of Explanation Queries as SAT Problems: Specific explanation queries are translated into logical constraints that are added to the encoded model and video data. For example, a ”Why?” query might involve finding a minimal set of input features or internal activations that are necessary for the observed prediction. A ”Why not?” query could involve finding minimal changes to the input that would lead to a different prediction.
@@
    SAT Solving and Explanation Extraction: A SAT solver is then used to find a satisfying assignment to the resulting set of logical clauses. The solution provided by the SAT solver is interpreted back in the context of the original model and video data to yield a formal explanation.
This framework aims to provide explanations that are not only interpretable but also formally verifiable, addressing the lack of rigor in many existing XAI techniques for deep learning in video understanding.
@@
Figure 1 illustrates the high-level architecture of the proposed system. It comprises a pre-trained deep learning model for video understanding, an encoding module that translates the model and the input video into a SAT problem, a query formulation module that defines the explanation query as logical constraints, a SAT solver, and an explanation extraction module that interprets the solver’s output.